OpenStudy (anonymous):

CALCULATE mass of 2.2 x 10^20 Ag atom. the answer is 0.0432 g but i keep getting 0.0394 g. Somebody pls help explain this to me!

OpenStudy (anonymous):

2.2x10^20/6.023x10^23 then multiply by mass of Ag

OpenStudy (anonymous):

I got 0.0394 as well

sam (.sam.):

using ratio method, \[2.2 \times 10^{20} \text{Ag atom} \times \frac{1~mol}{6.02 \times 10^{23}atom} \times \frac{107.9g}{1mol}=0.039g~Ag\] Your answer is correct

OpenStudy (anonymous):

thank you!